OpenStudy (appleduardo):

whats the derivative of this function? fx'=100/( x^(2) + y^(2) + z^(2) )

OpenStudy (appleduardo):

\[fx=\frac{ 100 }{ x ^{2} + y ^{2} +z ^{2}}\]

OpenStudy (loser66):



zepdrix (zepdrix):

\[\Large\rm f=100(x^2+y^2+z^2)^{-1}\]\[\Large\rm f_x=-100(x^2+y^2+z^2)^{-2} \color{royalblue}{(x^2+y^2+z^2)_x}\]\[\Large\rm f_x=-100(x^2+y^2+z^2)^{-2} (2x)\] Confused by any of that? The blue part was due to the chain rule.